Last year, Melissa had to $30000 invest. She invested some of it in an account that paid 6% simple interest per year, and she invested the rest in an account that paid 9% simple interest per year. After one year, she received a total of $1980 in interest. How much did she invest in each account?
Answer by mananth(16946)

Part I 6.00% per annum ------------- Amount invested =x
Part II 9.00% per annum ------------ Amount invested = y
30000
Interest----- 1980.00

Part I 6.00% per annum ---x
Part II 9.00% per annum ---y
Total investment
x + 1 y= 30000 -------------1
Interest on both investments
6.00% x + 9.00% y= 1980
Multiply by 100
6 x + 9 y= 198000.00 --------2
Multiply (1) by -6
we get
-6 x -6 y= -180000.00
Add this to (2)
0 x 3 y= 18000
divide by 3
y = 6000
Part I 6.00% $ 24000
Part II 9.00% $ 6000

CHECK
24000 --------- 6.00% ------- 1440.00
6000 ------------- 9.00% ------- 540.00
Total -------------------- 1980.00

m.ananth@hotmail.ca